ROTATIONAL SPECTRA OF ADDUCTS OF FORMALDEHYDE WITH FREONS

\underline{GOU QIAN}, GANG FENG, LUCA EVANGELISTI, and W. CAMINATI, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy; MONTSER-RAT VALLEJO LÓPEZ, ALBERTO LESARRI, Departamento de Quimica Fisica y Quimica Inorganica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; EMILIO COCINERO, Departamento de Quimica Fisica, Facultad de Ciencia y Tecnologia, Universidad del Pais Vasco, E-48080 Bilbao, Spain.

The rotational spectra of three 1:1 complexes of formaldehyde (H_{2}CO) with freons, i.e. difluoromethane (CH_{2}F_{2}), fluorochloromethane (CH_{2}FCl) and trifluorochloromethane (CF_{3}Cl), have been observed and assigned using pulsed jet Fourier transform microwave technique. Several isotopologues (including some ^{13}C species) have been measured in natural abundance. The tunnelling splittings have been measured in the first two adducts with relative intensity 1:3, due to the internal rotation of the formaldehyde moity along its symmetry axis. The barriers to this motion have been estimated by using a flexible model. For the latter two complexes, each of transition displays the hyperfine structures due to the quadrupolar effects of ^{35}Cl (^{37}Cl) nucleus. The dissociation energy has been estimated within the pseudo-diatomic approximation for all three complexes.